OpenStudy (anonymous):

what is the LCM of (x+3)and (x^2-9)

OpenStudy (anonymous):

Factorize the (x^2 - 9) first..

Parth (parthkohli):

\(x^2 - 9 = (x + 3)(x - 3)\)

OpenStudy (anonymous):

\[(a^2 - b^2) = (a+b)(a-b)\]

OpenStudy (anonymous):

(x+3)(x-3)

Parth (parthkohli):

\(x + 3\), \((x + 3)(x - 3)\). Now it's just oral.

OpenStudy (anonymous):

Now take the LCM.. One common factor and one non common factor will be the LCM..

Parth (parthkohli):

The common factor is \(x + 3\). The factors that are not common = LCM.

Parth (parthkohli):

I mean, common * uncommon

OpenStudy (anonymous):

(x+3)(x-3)??

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

THANKS!!

Parth (parthkohli):

That's right!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

:)

Parth (parthkohli):

\(x + 3 \times x - 3 = x^2 - 9\) \(x^2 - 9 \times 1 = x^2 - 9\) Confirmed

OpenStudy (anonymous):

lcm is the common tern in 2 equations (a2−b2)=(a+b)(a−b) x2−9=(x+3)(x−3) now common terms of (x+3)and (x^2-9) common terms of (x+3)and (x+3)(x−3) is the lcm